if a,b ∈ R with a < b, then there exists a rational number r ∈Qsuch that a < r <b.


1
Expert's answer
2021-01-19T04:01:15-0500

By Dedekind's definition of a real number, a real number "x" is characterised by all rational numbers that are smaller than this real number - "S_x=\{y\in\mathbb{Q}|y\leq x\}" . As "a<b", every rational number "p" such that "p\leq a" (which means "p\in S_a") satisfies also "p\leq b" by transitivity of an order (which means that "p\in S_b") , so "S_a\subset S_b". But as "b\neq a", we should have "S_b\neq S_a", so there is "r\in S_b, r\notin S_a", which means "r" is a rational number such that "a<r\leq b". If "r\neq b, a<r<b" and we have found the number we have been searching for. If "r=b, b\in \mathbb{Q}". In this case we can apply the same reasoning as above to find a rational number "r'" such that "-b<r'\leq-a". This gives us either a rational number "r', -b<r'<-a" and thus "a<-r'<b", or "r'=-a" and thus "a\in\mathbb{Q}". But in the latter case "a,b" are both rational and thus "\frac{a+b}{2}" is a rational number such that "a<\frac{a+b}{2}<b". Therefore we have found such a rational number in all the possible cases.